in a treasury bond quote with a 1000 face value you find the bid is equalt to 100-24 and thenask is equal to 1---26 you could buy this bond for 1008.25

True ( 100 + 26 * 1/32) = 1008.25

Can someone explain how the equation to the left equates to 1008.25

And explain where 1/32nds comes from.

I know the anwser is True, but i'm not sure how to prove that?.. please explain, thank you
Answer #1

True.
( 100 + (26 * 1/32)) = 1008.125

26 * 1/32 = 0.8125

Thus, 100 + 0.8125 = 100.8125

Face value = 1000

100.8125 * 1000 / 100 = 1008.125

You could buy this bond for 1008.125


Normal fraction used for Treasury security prices is 1/32.